OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -3x + 4y = -19.5 -3x + y = -10.5

hero (hero):

Hint: Multiply either equation by -1, but only one of them

OpenStudy (anonymous):

-3x+4y=-19.5(1) 3x-4y=19.5

OpenStudy (jlastino):

Then write it down like this 3x-4y=19.5 -3x+y=-10.5 -------------- Just add like you always do :D. The "x" will be "eliminated"

OpenStudy (anonymous):

thank you!